OpenStudy (anonymous):

1 points Save Match each with the expression or equation which best describes it. Question 1 answers Term Read Answer Items Coefficient Read Answer Items Associative Property Read Answer Items Distributive Property Read Answer Items Commutative Property Read Answer Items Answer A. a(b + c) = ab + ac B. This is what the 5 in the term 5xy is called. C. (a + b) + c = a + (b + c) D. This expression has 4 of them: 3a + b - 5 + 8d. E. a + b = b + a

OpenStudy (mertsj):

A coefficient is a number in front of a variable. For example, in the expression x2-10x+25, the coefficient of the x2 is 1 and the coefficient of the x is -10. The third term, 25, is called a constant.

OpenStudy (mertsj):

In Algebra a term is either: * a single number, or * a variable, or * numbers and variables multiplied together.

OpenStudy (anonymous):

help with the others pls i have to match them

OpenStudy (anonymous):

and thank you

OpenStudy (mertsj):

Notice that the SUM is the same no matter what way you group the addends. In general, the associative property of addition can be written as: (a + b) + c = a + (b + c)

OpenStudy (anonymous):

anymore pls like all of em pls

OpenStudy (mertsj):

As generally used in math class, the distributive property is the ability of one operation to "distribute" over another operation contained inside a set of parenthesis. Most commonly, this refers to the property of multiplication distributing over addition or subtraction, such that x(a+b) = xa + xb.

OpenStudy (mertsj):

Commutative property is one of the basic properties of numbers. The word “commute” means “exchange” or “swap over”. Commutative property states that numbers can be added or multiplied in any order. That is: Commutative Property of Addition states that changing the order of addends does not change the sum. That is, a + b = b + a.

OpenStudy (mertsj):

Do you have a computer?

OpenStudy (anonymous):

im on it

OpenStudy (mertsj):

Is there some reason, then, that you could not type those terms into a search engine on your own instead of asking someone else to do it for you?

OpenStudy (anonymous):

because it came up with this